A ball is falling at 7.3 m/s (down). Assuming Earth’s acceleration of gravity is 9.81 m/s2 (down), what will the balls final velocity be if it continues to fall for another 3.3 sec?
1
Expert's answer
2020-05-18T10:24:55-0400
"v_f=v_i+gt=7.3+(9.81)(3.3)=39.7\frac{m}{s}"